OpenStudy (anonymous):

Solve for x: 2 over quantity x minus 2 plus 7 over quantity x squared minus 4 equals 5 over x

OpenStudy (psymon):

\[\frac{ 2 }{ x-2 }+\frac{ 7 }{ x ^{2}-4 }=\frac{ 5 }{ x } \]? This correct?